# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

### v2

@@ 6 Analysis
How many examples are needed?
In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 3, 5, and 10. Our analysis reveals no significant differences in performance across these variations, while 3 self-generated examples are sufficient to improve performance. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
$k=3$ $k=5$ $k=10$
Gemini Pro 54.6 54.1 54.7
Gemini Ultra 54.7 54.1 53.8
Gemma-2b-it (v1.1) 41.2 40.8 41.0
Gemma-7b-it (v1.1) 51.3 50.2 49.5
Mistral-7b-instruct (v0.2) 52.8 52.8 52.4
Llama-2-7b-chat 33.3 32.5 31.5
Llama-2-13b-chat 34.9 34.4 33.0
Llama-2-70b-chat 51.6 51.1 51.5
Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper.
@@
To ablate the second principle in DoubleDipper, namely the explicit identification of the supporting paragraphs before generating the answer, we prompt Gemini Pro and Gemini Ultra with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)). Gemini Pro achieves on average 36.2 points (-18.4) and Gemini Ultra achieves 54.0 (-0.7 points). This indicates that instructing small models (e.g., Gemini Pro) with explicit prediction of supporting paragraphs achieves similar results to large models (Gemini Ultra).
Do few-shot examples really “instruct” the model to comprehend the text or provide clues to the query?
To answer this question, we prompt Gemini Pro and Gemini Ultra with only the self-generated demonstrations without the input context. As expected, removing the input context leads to a huge performance drop, -31.7 for Gemini Pro and -22.8 for Gemini Ultra compared to DoubleDipper. This confirms that DoubleDipper indeed “teaches” the model how to comprehend the input.
Qualitative Analysis
We manually analyze 50 prompts, with few-shot demonstrations produced by Gemini Ultra. Our review confirms that 93.5% of these self-generated QAs are correct, meaning that the question is meaningful and the answer could be found in the corresponding paragraph.
@@ 7 Conclusion
We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input. By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context. Experimental results show that our approach ourperforms both the vanilla LLM and the traditional ICL in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
@@ Appendix A Prompts
Figure 5: Template prompt for the generation question-answer pairs.
Appendix B Identification of Supporting Passages
Table 7 presents the F1 results of the tested models for the supporting relevant passages identification. Without any demonstration (zero-shot), all open source models achieve a poor performance, while DoubleDipper significantly improves performance (+37.8 F1 for Llama2 70B, +38 F1 for Gemma 7B, +30.5 F1 for Mistral).
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Baseline - - - - - - - -
Zero-shot 54.5 58.2 47.1 75.9 46.6 56.3 58.3 39.4
DoubleDipper (Self) 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper (PaLM 2) 50.6 60.1 47.3 68.0 21.2 58.8 57.9 41.0
Gemini Ultra Baseline - - - - - - - -
Zero-shot 40.4 57.4 30.3 39.4 17.0 52.4 43.7 42.9
DoubleDipper (Self) 37.5 58.6 25.5 21.5 11.1 55.4 46.4 44.3
DoubleDipper (PaLM 2) 38.4 58.0 28.0 24.6 12.0 55.3 47.0 43.7
Gemma-2b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 23.3 10.4 43.1 37.5 13.5 20.1 13.0 25.3
DoubleDipper (PaLM 2) 25.6 9.1 48.7 48.7 15 18.7 14 25.3
Gemma-7b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 38.0 39.7 44.8 59.8 7.1 41.8 36.5 36.1
DoubleDipper (PaLM 2) 36.9 39.2 45.2 58.6 7.3 40.1 32.8 35
Mistral-7b-instruct (v0.2) Baseline - - - - - - - -
Zero-shot 11.5 6.7 0.7 2.9 0.0 17.2 18.2 34.5
DoubleDipper (Self) 42.0 48.2 66.2 60.6 5.8 41.7 29.3 42.2
DoubleDipper (PaLM 2) 41.5 48.3 70.0 58.6 6.8 40.0 26.0 40.6
Llama-2-7b-chat Baseline - - - - - - - -
Zero-shot 0.8 0.2 0.0 0.0 3.2 0.0 0.0 2.1
DoubleDipper (Self) 19.4 17.5 27.7 20.2 5.7 21.7 19.7 23.3
DoubleDipper (PaLM 2) 19.6 15.8 29.3 21.6 5.7 22.2 20.7 22.0
Llama-2-13b-chat Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 28.8 16.5 51.0 32.0 18.4 29.6 22.2 31.7
DoubleDipper (PaLM 2) 30.3 13.4 51.6 30.5 15.6 36.4 27.3 36.9
Table 7: Performance (F1) of supporting paragraph(s) prediction. The baseline does not predict supporting paragraphs. DoubleDipper provides a significant performance boost for all open source models, whose performance is close to 0 (except from Mistral) in the zero shot setting.

### v3

@@ 6 Ablation Studies
To ablate the second principle in DoubleDipper, which is the explicit identification of the supporting paragraphs before generating the answer, we prompt the open source models with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)). For all models, the average QA performance of DoubleDipper without the evidence is lower than DoubleDipper with evidence identification. When averaging results across models and datasets, we report a substantial F1 drop from 54.8 to 46.6. Full results are available in Appendix C.3 in Table 9.
Qualitative analysis: Correctness of the generated QA pairs
We manually analyze 150 QAs generated by Gemma 2B as demonstrations. Our review confirms that 93.5% of these self-generated QAs are correct, meaning that the question is meaningful and the answer could be found in the corresponding paragraph.
@@ 7 Conclusion
We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input. By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context. Experimental results show that our approach substantially ourperforms the baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
@@ Appendix A Prompts
Figure 5: Template prompt for generating the QA pairs in DoubleDipper.
Appendix B Lost-in-the-middle
Figure 6 shows the QA accuracy of the models Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro on our subset of the “Lost-in-the-middle” dataset.
Figure 6: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
Appendix C Analysis
C.1 Impact of the Number of Demonstrations in DoubleDipper
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemini Nano $k=1$ 60.03 37.68 85.20 88.20 56.40 62.50 60.68 29.56
$k=3$ 62.12 40.55 86.60 95.40 56.20 65.12 60.52 30.44
$k=5$ 62.25 41.79 87.00 95.60 55.20 65.05 60.56 30.52
$k=10$ 62.33 43.16 86.00 96.20 55.20 65.37 60.44 29.96
Gemini Flash $k=1$ 65.34 48.83 84.80 90.80 66.40 60.17 65.16 41.19
$k=3$ 66.11 48.03 85.80 95.00 68.60 60.58 65.04 39.71
$k=5$ 65.87 47.49 87.20 95.00 66.60 61.13 64.20 39.48
$k=10$ 66.08 46.13 86.20 95.20 69.00 62.03 63.80 40.18
Gemma 2B (v2) $k=1$ 47.05 24.05 73.60 77.00 50.20 41.32 47.04 16.13
$k=3$ 49.48 23.66 85.80 81.60 50.00 39.85 46.68 18.77
$k=5$ 49.59 26.70 85.40 80.40 48.40 40.34 46.56 19.30
$k=10$ 49.92 26.43 85.80 81.40 49.00 40.91 47.68 18.26
Gemma 9B (v2) $k=1$ 58.77 37.58 81.00 87.60 51.60 58.46 63.44 31.68
$k=3$ 61.20 41.70 84.00 95.00 51.40 61.21 61.80 33.31
$k=5$ 61.40 44.42 84.20 94.80 50.00 62.99 60.76 32.60
$k=10$ 61.34 43.78 84.40 95.80 50.40 63.01 59.96 32.05
Llama 3.1 8B $k=1$ 57.74 38.20 85.40 84.80 51.80 57.50 60.96 25.49
$k=3$ 59.86 38.75 91.20 90.60 51.00 59.29 58.12 30.09
$k=5$ 60.58 41.56 89.80 91.40 51.00 61.31 57.68 31.28
$k=10$ 61.41 44.68 89.00 91.00 51.00 63.88 56.36 33.98
Mistral 7B $k=1$ 48.91 28.86 67.60 77.40 50.00 40.62 62.00 15.91
$k=3$ 51.03 28.65 68.40 88.80 50.60 43.44 60.68 16.66
$k=5$ 51.12 29.86 69.20 89.00 49.40 43.83 60.36 16.23
$k=10$ 51.44 30.14 68.20 89.40 49.00 46.49 59.96 16.91
Mixtral 7x8B $k=1$ 49.26 19.83 86.00 74.80 49.40 31.73 69.00 14.07
$k=3$ 52.22 22.31 91.80 86.00 47.80 35.10 66.56 15.97
$k=5$ 51.65 23.55 91.40 82.40 47.20 34.64 64.92 17.46
$k=10$ 52.18 23.49 91.80 84.40 47.00 36.23 64.00 18.32
Table 7: Performance of DoubleDipper on our evaluation set with various numbers of self-generated few shot demonstrations ($k$).
Table 7 presents the results of DoubleDipper with 1, 3 (main experiment in the paper), 5 and 10 generated demonstrations. For all these experiments, the demonstrations were generated by Gemma 2 2B.
Figure 7 shows the QA accuracy of DoubleDipper on “Lost” according to the position of the relevant passage for each $k\in\{1,3,5,10\}$.
Figure 7: Performance (accuracy) of several models with DoubleDipper according to the number of self-generated demonstrations in the prompt.
C.2 Impact of the few-shot generator
Table 8 presents the detailed QA performance of all models with different models for generating DoubleDipper’s demonstrations. As mentioned in the paper (Section 6), generating the demonstrations with the best model (ie. Gemini Pro) achieves the best performance overall.
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemini Pro Gemma 2B 70.4 46.8 97.4 99.0 79.6 60.9 72.4 36.4
Self 71.6 49.6 98.2 99.6 78.8 62.5 72.5 40.1
Gemini Flash Gemma 2B 66.1 48.0 85.8 95.0 68.6 60.6 65.0 39.7
Self 67.5 49.5 88.6 96.8 65.6 64.3 65.2 42.5
Gemini Pro 68.1 50.3 90.4 96.0 69.0 64.8 64.8 41.6
Gemini Nano Gemma 2B 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
Self 61.7 39.8 85.8 95.2 56.2 64.3 60.7 30.0
Gemini Pro 62.8 40.4 87.2 99.0 54.4 66.8 60.4 31.7
Gemma 2 2B Self 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
Gemini Pro 51.0 25.2 83.8 93.4 48.0 41.7 46.2 18.4
Gemma 2 9B Gemma 2B 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
Self 62.0 44.8 82.4 97.0 50.6 62.9 61.5 34.9
Gemini Pro 63.5 46.2 86.4 99.2 52.2 64.4 61.4 34.8
Llama 3.1 8B Gemma 2B 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
Self 60.1 41.3 88.6 89.8 50.8 61.5 57.8 31.0
Gemini Pro 61.5 41.0 89.8 95.6 51.6 63.5 57.5 31.7
Mistral 7B (v0.3) Gemma 2B 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
Self 49.8 25.2 68.8 93.8 49.6 36.5 60.6 14.3
Gemini Pro 52.6 31.4 67.2 96.4 48.2 45.7 62.1 17.0
Mixtral 7x8B (v0.1) Gemma 2B 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
Self 49.8 16.0 91.8 86.0 48.8 25.5 67.8 12.8
Gemini Pro 54.4 20.8 94.8 96.2 49.0 36.3 66.6 16.9
Table 8: Performance of DoubleDipper according to the model for generating the demonstrations (Gemma 2B, Self or Gemini Pro).
C.3 Impact of the identification of supporting paragraphs in the QA generation
Table 9 compares the performance of DoubleDipper to DoubleDipper without evidence identification.
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemma 2B DoubleDipper (QA only) 46.4 16.0 84.6 79.2 49.6 31.1 49.5 14.7
DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
Gemma 9B DoubleDipper (QA only) 55.1 22.5 91.0 96.6 52.0 42.9 61.0 20.0
DoubleDipper 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
Mistral 7B (v0.3) DoubleDipper (QA only) 49.1 21.8 73.4 89.8 48.6 38.6 60.3 11.5
DoubleDipper 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
Mixtral 8x7B (v0.1) DoubleDipper (QA only) 49.5 17.0 91.4 86.8 50.8 22.9 65.8 11.9
DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
Llama 3.1 8B DoubleDipper (QA only) 32.7 25.0 9.2 41.4 46.6 33.5 56.7 16.6
DoubleDipper 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
Table 9: Performance of DoubleDipper and DoubleDipper without instructing the models to retrieve the evidence (QA only) on the QA datasets.